OpenStudy (anonymous):

A train of length 100m is moving at a speed of 50km/h. A horse is running alongside the train at a speed of 56km/h. How long will it take the horse to overtake the train?

OpenStudy (anonymous):

distance = 100 m = 100/1000 = 0.1 km speed of train = 50 km/hr speed of horse = 56 km/hr resultant speed = (56 – 50) 6 km/hr time = distance/speed time = 0.1/6 = 1/60 hr = 1 min distance = speed × time time = 1 hr = 3600 s. = 300000 × 3600 = 1080000000 = 1.08 × 10^9 km ---

OpenStudy (solomonzelman):

no, the train is not moving and horse is going 6km/h.

OpenStudy (solomonzelman):

\[\frac{6km}{60\min}=\frac{100m}{x}\]

OpenStudy (solomonzelman):

\[---> \frac{1km}{10}=\frac{.1km}{x}\] srocc multiply to solve for x.

OpenStudy (solomonzelman):

cross multiply, my bad,.

OpenStudy (anonymous):

thank you !

OpenStudy (solomonzelman):

Anytime!\[(•_‿•)\]